An economy's aggregate demand curve (the relationship between short-run equilibrium output and inflation) is described by the equation:

Y = 11,000 - 17,000π, where π is the inflation rate.

Initially, the inflation rate is 3 percent or π = 0.03. Potential output Yp equals 10,660.

Note: Keep as much precision as possible during your calculations. Your final answer for inflation should be accurate to at least two decimal places and output should be accurate to the nearest whole number.

a) Find inflation and output in short-run equilibrium.
b) Find inflation and output in long-run equilibrium
1
Expert's answer
2020-11-17T12:44:23-0500

a) "Y=11000-17000\pi"

"\pi=0.03" or "3 \%"

Then "Y=11000-17000\times0.03=10490"

b) Potential output "Y_{p}=10660"

"10660=11000-17000\pi"

"-340=-17000\pi"

"\pi=0.02" or "2\%"

In the long run the equilibrium output level is the level of potential output that is 10660.